What is the probability that the 36 tires will have an

average of less than 16,000 miles until the tires begin

to wear out


1
Expert's answer
2022-05-16T17:52:33-0400

We have that: "X\sim N(\mu, \sigma^2/n)"

"\mu=16800"

"\sigma=3300"

"n=36"

"P(\bar X<16000)=P(Z<\dfrac{X-\mu}{\sigma/\sqrt{n}})""=P(Z<\dfrac{16000-16800}{3300/\sqrt{36}})=P(Z<-1.454545)"

"\approx0.0729"